OpenStudy (anonymous):

f(x)=sin ( 2x + pi/3 ) 1)a) préciser la peridoe de f b) montrer que delta : x = 7pi/12 est un awe de symétrie de Cf c) on déduire un domaine d'étude de f 2) etudier les variations de f et tracer la courbe dans un repere orthnormé 3) soit g(x)=f(x-pi/2) Tracer Cg a partir de Cf #Michele_Laino @Michele_Laino

OpenStudy (michele_laino):

By definition of a period T of a periodic function f, we have: f(x+T)=f(x) Now, applying that condition to our function, we get: \[\sin \left( {2\left( {x + T} \right) + \frac{\pi }{3}} \right) = \sin \left( {2x + \frac{\pi }{3}} \right)\] Developing the necessary computation, we get the subsequent condition: \[\Large \cos \left( {2T} \right) = 1\], please check that formula. So what is T?

OpenStudy (anonymous):

f (x) = sin (2x + pi / 3) 1) a) specify the reporting period of f b) show that delta x = 7ft / 12 is an axe symmetry Cf c) deriving a field of study of f 2) study the variations of f and graph it in a repere orthnormé 3) be g (x) = f (x-pi / 2) Tracer Cg from Cf

OpenStudy (michele_laino):

the above quantity T, answers to question a)

OpenStudy (anonymous):

yes thank you

OpenStudy (michele_laino):

please, what is Cf?

OpenStudy (anonymous):

la courbe de f

OpenStudy (michele_laino):

the graph of f(x)

OpenStudy (anonymous):

yes

OpenStudy (michele_laino):

let's suppose that x=a, is a symmetry axis, then we can write: f(x)=f(2a-x) substituting our function, we get: \[\Large \sin \left( {2x + \frac{\pi }{3}} \right) = \sin \left( {2\left( {2a - x} \right) + \frac{\pi }{3}} \right)\] then, starting from that condition, what value do you get for a?

OpenStudy (anonymous):

7pi/12

OpenStudy (michele_laino):

are you sure?

OpenStudy (michele_laino):

when we have the subsequent equality: sin x = sin y

OpenStudy (michele_laino):

do you know that answer?

OpenStudy (anonymous):

no xD to be honest i never passed math exam this year

OpenStudy (michele_laino):

the answer is: when the subsequent condition holds: x= pi- y

OpenStudy (michele_laino):

\[\Large x = \pi - y\]

OpenStudy (michele_laino):

so we have to solve this equation for a: \[\Large 2x + \frac{\pi }{3} = \pi - \left\{ {2\left( {2a - x} \right) + \frac{\pi }{3}} \right\}\] what is a?

OpenStudy (anonymous):

what is a ? the professor said x= 7pi/12

OpenStudy (michele_laino):

There I keep in mind the subsequent drawing: